OpenStudy (anonymous):

Whats an example of how rational exponets help you to multiply two radicals with different indices ?

OpenStudy (amistre64):

you can multiply like bases and add the exponents,,,,,

OpenStudy (amistre64):

\[\sqrt{5} * \sqrt[3]{5}\]

OpenStudy (amistre64):

\[5^{1/2} * 5^{1/3} = 5^{1/2 + 1/3}\]